# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

### v1

@@ C.3 Proof for Matching Regret Upper Bound in Theorem 1
Through interactions with the user, the learner $\mathcal{L}$ endeavors to minimize the dueling regret, as defined below.
$\text{Reg}^{\text{DB}}_{T}=\mathbb{E}\left(\sum^{T}_{t=1}\sigma(f(a_{t})-f(a^{*}))+\sigma(f(a^{\prime}_{t})-f(a^{*}))-1\right).$ (16)
Our minimization objective in (16) aligns with the regret definition in Yue and Joachims (2009). Under Assumptions C.1, C.2, and C.3, the regret defined in (16) is equivalent to the regret defined in Kumagai (2017). Assumption C.2 suggests the existence of a unique minimizer $a^{*}$ for the cost function $f$ due to its strict convexity. Assumptions C.2 and C.4 imply that the diameter $R:=\sup_{a,a^{\prime}\in\mathcal{A}}|a-a^{\prime}|$ and $B$ are finite. Assumption C.3 implies the existence of positive constants $l_{0}$, $L_{0}$, $B_{2}$, and $L_{2}$ such that the first derivative $\sigma^{\prime}$ of the link function is bounded, satisfying $l_{0}\leq\sigma^{\prime}\leq L_{0}$ on $[-B,B]$. Furthermore, the second derivative $\sigma^{{}^{\prime\prime}}$ is bounded above by $B_{2}$ and is $L_{2}$-Lipschitz on $[-B,B]$. Lemma C.7 establishes the relationship between $\text{Reg}^{\text{FO}}_{T}$ and $\text{Reg}^{\text{DB}}_{T}$.
Lemma C.7 (Lemma 12 in Kumagai (2017)).
$\frac{\text{Reg}^{\text{DB}}_{T}}{L_{0}}\leq\text{Reg}^{\text{FO}}_{T}\leq\frac{\text{Reg}^{\text{DB}}_{T}}{l_{0}}.$ (17)
C.3.1 Noisy Comparison-based Stochastic Mirror Descent
Input: Learning rate $\eta_{\rho}$, $\nu$-self-concordant function $\mathcal{R}$, time horizon $T$, tuning parameters $\lambda$, $\mu$.
1 Initialize $a_{1}=\arg\min_{a\in A}\mathcal{R}(a)$
2 for $t\in[T]$ do
3       Update $\mathcal{R}_{t}(a)=\mathcal{R}(a)+\frac{\lambda\eta_{\rho}}{2}\sum^{t}_{i=1}\|a-a_{i}\|^{2}+\mu\|a\|^{2}$
4       Pick a unit vector $u_{t}$ uniformly at random
5       Receive the corrupted feedback ^$\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})$, for $a_{t}$ and $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-1/2}u_{t}$
6       Compute the corrupted gradient: $\hat{g}_{t}=\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})d\nabla^{2}\mathcal{R}_{t}(a_{t})^{1/2}u_{t}$
Output : $a_{T+1}$
Algorithm 3 Noisy Comparison-based Stochastic Mirror Descent (NC-SMD)
Remark C.1.
Note that $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}x$ for $x\in\mathbb{B}$ is included in $\mathcal{A}$ due to the properties of the Dikin ellipsoid.
Now we begin to prove the matching regret upper bound in Theorem 1.
Proposition C.1 (Full Statement of Matching Regret Upper Bound in Theorem 1).
Let the number of rounds $T$ be fixed and $T>\sqrt{2LL_{0}R}$. Consider a corruption induced by $\rho$-imperfect user (Def. 1) with known $\rho$. Under Assumptions C.2, C.3, and C.4, choose $\lambda\leq l_{0}\alpha/2$, $\mu\geq(L^{3}_{0}L_{2}/\lambda)^{2}$, and set the learning rate $\eta_{\rho}:=\frac{\sqrt{\log T}}{dT^{\max\{1/2,\rho\}}}$ for Algorithm 3. We have we have:
$\text{Reg}^{\text{DB}}_{T}\leq O\left(\max\left\{d\sqrt{T\log T},dT^{\rho}\log T\right\}\right).$
Proof  (Proof for Proposition C.1). The formal proof of Proposition C.1 relies on Lemmas C.8, C.9, and C.10. Lemma C.8 establishes a connection between the gradient estimation corrupted by the bias term $b_{t}$ and the quantity $\bar{P}_{t}(a)$ (defined in Lemma C.8), which is the main component we aim to analyze in the regret definition (16). Lemma C.9 decomposes the regret into two parts: a regret of decision independent of $b_{t}$ and an error in observation summarizing the cumulative impact of $b_{t}$.
The optimization error can be explicitly upper bounded by $T$ and $\eta_{\rho}$ using techniques similar to those employed in Kumagai (2017). The upper bound of the estimation error is further established in Lemma C.10, wherein we utilize the $\alpha$-strong convexity of $f$ to establish a connection between the convergence of the $L_{2}$-norm of the action $\|a_{t}-a^{*}\|_{2}$ to itself.
Lemma C.8 (Gradient Estimation under Corruption).
Let $P_{t}(a)$ represent the likelihood of action $a_{t}$ being preferred over $a$, where $P_{t}(a):=\sigma(f(a)-f(a_{t}))$. Likewise, the corrupted version is defined as $\hat{P}_{t}(a)=\sigma(f(a)-f(a_{t})+c_{t}(a_{t},a^{\prime}_{t}))$. The smoothed version of $P_{t}$ over $\text{int}(\mathcal{A})$ is $\bar{P}_{t}(a):=\mathbb{E}_{x\in\mathbb{B}}\left[P_{t}(a+\nabla^{2}R_{t}(a_{t})^{-\frac{1}{2}}x)\right]$. $\mathcal{R}_{t}(a)$ is defined in Algorithm 3. Let $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}$, where $u_{t}$ is uniformly sampled from $\mathbb{S}$. Let $b_{t}:=d\left(P_{t}(a^{\prime}_{t})-\hat{P}_{t}(a^{\prime}_{t})\right)\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}$, we have
$\mathbb{E}\left(\hat{g}_{t}|a_{t}\right)=\nabla\bar{P}_{t}(a_{t})-\mathbb{E}\left(b_{t}|a_{t}\right).$
Lemma C.9 (Regret Decomposition).
Define $C_{0}:=(2\nu+4L_{0}\beta+4B_{2}L^{2}+LL_{0}\beta)/\lambda$, and let $a^{*}_{T}=\frac{1}{T}a_{1}+(1-\frac{1}{T})a^{*}$. By Assumption C.3, executing Algorithm 3, we obtain:
[…]
Lemma C.10 (Estimation Error Upper Bound).
Under Assumption C.2, and C.4, we have
$\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$ $\leq d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right\}+C_{2}dC+2d\sqrt{\lambda\eta_{\rho}T},$
where we have $\lambda^{*}_{\mathcal{R}}:=\sup_{a\in\mathcal{A}}\lambda_{\max}\left(\nabla^{2}\mathcal{R}(a)\right)$, and $m_{t},n_{t},C_{2}$ defined as
$m_{t}:=\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right),n_{t}:=\min\left(2R,\sqrt{\frac{2}{\alpha}\left(f(a_{t})-f(a^{*})\right)}\right),C_{2}:=2L_{\sigma}R\sqrt{\lambda^{*}_{R}+2\mu}.$
Now, we present the proof of Proposition C.1. The key technical innovation lies in realizing that the upper bound of $\text{Reg}^{\text{DB}}_{T}$ contains itself. This implies that we can initially obtain a rough upper bound and then refine it until it becomes tight. The proof follows a mathematical induction argument. For the ease of notation, we reparameterize $C=C_{\kappa}T^{\rho}$. In the following proof, we consider $\rho>0.5$ to make $T^{\rho}=\max\{\sqrt{T},T^{\rho}\}$. If $\rho\leq 0.5$, the following proof can be easily modified to make upper bounds involving $T^{\rho}$ as $\sqrt{T}$.
In the first step, we employ a rough upper bound, specifically $\mathbb{E}\left(n_{t}\right)\leq 2R$. This yields $\text{Reg}^{DB}_{T}\leq O(dT^{\frac{\rho+1}{2}})$. Since $\text{Reg}^{DB}_{T}$ and $\text{Reg}^{\text{FO}}_{T}$ share the same order (see Equation (17)), it implies that $\sum^{T}_{t=1}\mathbb{E}\left(n_{t}\right)$ must be sublinear as well. Therefore, in the second step, we can use Abel’s Summation Equation (Lemma C.16) to refine the upper bound of $\mathbb{E}\left(\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right)$.
In particular, we rigorously prove the following claim:
The Induction Claim: Given the number of iteration $T$. $T>\sqrt{2LL_{0}R}$. At k-th step, we have $\text{Reg}^{\text{DB}}_{T}\leq 144dK\sqrt{\log T}T^{\rho+\frac{\frac{3}{2}-\rho}{2^{k+1}}}$, where $K:=\max\left\{C_{0},\frac{C_{2}C_{\kappa}}{\sqrt{\log T}},L^{2}_{\sigma}C^{2}_{\kappa},2\sqrt{\lambda}RC_{\kappa}L_{\sigma}\right\}$.
The Base Case: When $k=1$, let $C_{3}=C_{2}dC+2d\sqrt{\lambda\eta_{\rho}T}$, we have
$\text{Reg}^{\text{DB}}_{T}$ $\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+2C_{3}+2d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left(\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right).$ (18)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+2C_{3}+4dR\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}m_{t}\right\}$ (19)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+2C_{3}+4dR\sqrt{\lambda\eta_{\rho}T}\left(\sum^{T}_{t=1}m_{t}\right)$ (20)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+2C_{3}+4dL_{\sigma}R\sqrt{\lambda\eta_{\rho}}C_{\kappa}T^{\rho}.$ (21)
We get Equation (18) according to Lemma C.9 and C.10. We get Equation (19) because of $\mathbb{E}\left(n_{t}\right)\leq 2R$. We get Equation (20) by Abel’s Summation Equation (see Lemma C.16). We get get Equation (21) because of corruption budget. Choosing the learning rate $\eta_{\rho}$ according to Proposition C.1, we obtain $\text{Reg}^{\text{DB}}_{T}\leq 144\sqrt{\log T}KdT^{\frac{\rho+1}{2}}$. This confirms the validity of the claim when $k=1$.
The Induction Argument: Let’s assume that the claim holds true for a general step $k$. In the following, we will demonstrate that the claim also holds true for step $k+1$. Because of Equation (17) and the induction claim, we obtain $\text{Reg}^{\text{FO}}_{T}\leq\frac{144}{l_{0}}dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k+1}}}.$ This suggests that
$\sum^{T}_{t=1}\mathbb{E}\left(f(a_{t})-f(a^{*})\right)\leq\frac{144}{l_{0}}dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k+1}}}.$
By Jensen’s inequality, we have
$\mathbb{E}(n_{t})=\min\left\{2,\sqrt{\frac{2}{\alpha}}\mathbb{E}\left(\sqrt{f(a_{t})-f(a^{*})}\right)\right\}\leq\sqrt{\frac{2}{\alpha}}\sqrt{\mathbb{E}\left(f(a_{t})-f(a^{*})\right)}.$
Therefore we have
$\mathbb{E}\left(\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right)\leq\sqrt{\frac{2}{\alpha}}\sum^{T}_{t=1}\sqrt{t}m_{t}\sqrt{\mathbb{E}\left(f(a_{t})-f(a^{*})\right)}.$
By the definition of monotone adversary, we have $|c_{t}(a_{t},a^{\prime}_{t})|$ is upper bounded by $c_{k}t^{-1+\rho}$.
Lemma C.11.
Therefore, using Lemma C.11 (details of proof is listed at Appendix C.3.5), we can upper bound $\sqrt{\lambda}\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}$ by
$\sqrt{\lambda}\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\leq 60L_{\sigma}\sqrt{\frac{2\lambda}{\alpha l_{0}}}(\log T)^{0.25}C_{k}\sqrt{dK}T^{\frac{3}{2}\rho+\frac{3/2-\rho}{2^{k+2}}}$
Since $\lambda\leq\alpha l_{0}/2$, choosing $\eta_{\rho}=\frac{1}{2d}\frac{\sqrt{\log T}}{T^{\rho}}$, we have
$\text{Reg}^{\text{DB}}_{T}$ $\leq 2d\left(C_{0}+\frac{C_{2}}{\sqrt{\log T}}+60C_{\kappa}\sqrt{K}T^{\frac{3/2-\rho}{2^{k+2}}}\right)\sqrt{\log T}T^{\rho}+12Kd\sqrt{\log T}T^{\rho}$
$\leq 144dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k+2}}}.$
$\text{Reg}^{\text{DB}}_{T}\leq O\left(d\sqrt{\log T}T^{\rho}\right),$
which completes the proof when $\rho\in[0.5,1]$.
C.3.2 Proof for Lemma C.8: Gradient Estimation Under Corruption
Proof  (Proof for Lemma C.8). If we do not have corruption (i.e. the probability for observing noisy comparative feedback $\mathcal{F}(a,a^{\prime})=1$ is based on true cost difference $f(a^{\prime})-f(a)$), then the uncorrupted gradient $g_{t}:=\mathcal{F}(a^{\prime}_{t},a_{t})d\nabla^{2}\mathcal{R}_{t}(a_{t})^{1/2}u_{t}$ should be an unbiased estimate of $\nabla\bar{P}_{t}(a_{t})$, i.e.
@@ 7       Set $a_{t+1}=\nabla\mathcal{R}_{t}^{-1}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t})$
The proof is similar to Lemma 2.1 in Flaxman et al. (2005). Using the Law of total expectation, we have
$\mathbb{E}(g_{t}|a_{t})$ $=\mathbb{E}_{u_{t}}(\mathbb{E}(g_{t}|a_{t},u_{t}))$
@@
C.3.3 Proof for Lemma C.9: Regret Decomposition
Proof  (Proof for Lemma C.9). The cornerstone of the analysis below is to separate the impact of $b_{t}$ from the total regret.
$\text{Reg}^{\text{DB}}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T})\right]+\frac{LL_{0}\beta}{\lambda\eta}+2LL_{0}R$
@@
$+\frac{LL_{0}\beta}{\lambda\eta}+2LL_{0}R$
$\leq 2\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T}))\right\}+\frac{4L_{0}\beta+4B_{2}L^{2}+LL_{0}\beta}{\lambda\eta_{\rho}}\log T+2LL_{0}R.$
We get the first inequality by using Lemma C.12. We get the last inequality by because
$\bar{P}_{t}(a)-P_{t}(a)\leq\frac{L_{0}\beta+B_{2}L^{2}}{2}\|\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}\|^{2}\leq\frac{L_{0}\beta+B_{2}L^{2}}{\lambda\eta_{\rho}t},$
@@
$\mathbb{E}\left\{\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T}))\right\}$ $\leq\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T}))\right\}+2\frac{L_{0}\beta+B_{2}L^{2}}{\lambda\eta_{\rho}}\log T.$
Lemma C.12 (Lemma 5 in Kumagai (2017)).
$\text{Reg}^{\text{DB}}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T})\right]+\frac{LL_{0}\beta}{\lambda\eta}+2LL_{0}R.$
@@
$\mathbb{E}\{\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T})\}$ $\leq\mathbb{E}\left[\nabla\hat{P}^{\top}_{t}(a_{t})(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[g_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[(\hat{g}_{t}+b_{t})^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]+\mathbb{E}\{b^{\top}_{t}(a_{t}-a^{*}_{T})\},$
where the first and second equality is resulted from Lemma C.8. Using the definition of $a_{t+1}$ in Algorithm 3, we have
@@
$\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$ $=\frac{1}{\eta_{\rho}}\mathbb{E}\left[(\nabla\mathcal{R}_{t}(a_{t+1})-\nabla\mathcal{R}_{t}(a_{t}))^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha\eta_{\rho}}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\frac{1}{\eta_{\rho}}\mathbb{E}\left[D_{\mathcal{R}_{t}}(a^{*}_{T},a_{t})+D_{\mathcal{R}}(a_{t},a_{t+1})-D_{\mathcal{R}_{t}}(a^{*}_{T},a_{t+1})-\frac{l_{0}\alpha\eta_{\rho}}{4}\|a_{t}-a^{*}_{T}\|^{2}\right].$
@@
Then we can get Equation (22) by using Lemma C.13, C.14, C.15.
@@
Lemma C.13 (Lemma 10 in Kumagai (2017)).
@@
$\sum^{T}_{t=1}\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]\leq\frac{1}{\eta_{\rho}}\left(\mathcal{R}(a^{*}_{T})-\mathcal{R}(a_{1})+\mathbb{E}\left[\sum^{T}_{t=1}D_{\mathcal{R}^{*}_{t}}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t},\nabla\mathcal{R}_{t}(a_{t}))\right]\right).$
Lemma C.14 (Lemma 11 in Kumagai (2017)).

### v2

@@ B Proof for Theorem 2
Additionally, we remind the reader of the standard assumptions (Yue and Joachims, 2009; Kumagai, 2017) used in our analysis.
  1. 1.
    the action set $\mathcal{A}$ is a convex compact set that contains the origin, has non-empty interior, and is contained in a $d$-dimensional ball of radius $R$;
  2. 2.
    the utility function $\mu:\mathcal{A}\rightarrow\mathbb{R}$ is strongly concave and twice-differentiable. The following constants are useful for our algorithm analysis: $\mu$ is $L^{\mu}$-Lipschitz, $\alpha$-strongly concave, $\kappa$-smooth, bounded $R^{\mu}:=\sup_{a,a^{\prime}}\mu(a)-\mu(a^{\prime})$, and $a^{*}:=\arg\max_{a\in\mathcal{A}}\mu(a)$ is the unique optimal within $\mathcal{A}$;
  3. 3.
    the link function $\sigma:\mathbb{R}\rightarrow[0,1]$ is smooth, rotation-symmetric (i.e. $\sigma(x)=1-\sigma(-x)$), and concave for any $x\geq 0$. For the ease of analysis: let $l^{\sigma}_{1}$ [resp. $L^{\sigma}_{1}$] denote the lower [resp. upper] bound of the first-order derivative of $\sigma$. $\sigma$ is $L^{\sigma}$-Lipschitz and its second-order derivative $\sigma^{\prime\prime}$ is $L^{\sigma}_{2}$-Lipschitz and upper bounded by $R^{\sigma}_{2}$.
Under these assumption, Lemma 15 implies $\text{Reg}_{T}$ defined in (1) and $\text{Reg}^{\text{FO}}_{T}$ defined in (3) has the same order.
Lemma 15 (Lemma 12 in Kumagai (2017)).
With $\text{Reg}^{\text{FO}}_{T}$ defined in (3), we have
$\frac{\text{Reg}_{T}}{L^{\sigma}_{1}}\leq\text{Reg}^{\text{FO}}_{T}\leq\frac{\text{Reg}_{T}}{l^{\sigma}_{1}}.$ (14)
In the following, we show the proof. The main proof can be divided into four key steps. Step 1 is natural, which quantifies the bias in gradient estimation by RoSMID due imperfect feedback from corrupted utilities. Building on this, Step 2 develops a regret decomposition lemma for dueling bandits under corruption, decomposing the regret into two components: regret from sub-optimal decisions and regret from feedback error. Techniques from previous analyses of continuous dueling bandits handle the sub-optimal decision regret, thus in Step 3, we focus on deriving new techniques to bound the regret from feedback error, a unique challenge in our setting. We identify a tight upper bound for this error. Finally, Step 4, the only step relying on the decaying corruption level in $\rho$-imperfect human feedback, uses this assumption, along with the carefully chosen learning rate $\eta_{\rho}$ and an iterative regret refinement process, to ensure the tightness of the analysis in Step 3, thus completing the proof.
Step 1: quantifying bias of gradient estimation in $\rho$-LIHF
Our proof starts from quantifying the bias of the gradient $\hat{g}_{t}$ caused by $\rho$-imperfect human feedback, as shown in the following Lemma 2.
Lemma (Lemma 2 restated).
The gradient $\hat{g}_{t}$ in Line 6 of Algorithm 1 satisfies
$\mathbb{E}\left(\hat{g}_{t}|a_{t}\right)=\nabla|_{a=a_{t}}\bar{P}_{t}(a)-\mathbb{E}\left(b_{t}|a_{t}\right).$
where $\bar{P}_{t}(a):=\mathbb{E}_{x\in\mathbb{B}}\left[\mathbb{P}(a_{t}\succ a+\nabla^{2}R_{t}(a_{t})^{-\frac{1}{2}}x)\right]$ (with $\mathbb{B}$ as the unit ball) is the smoothed probability that $a_{t}$ is preferred over any action $a$, and $b_{t}=d\left(\mathbb{P}(a_{t}\succ a^{\prime}_{t})-\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})\right)\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}$.
Proof  If we do not have corruption (i.e. the probability for observing noisy comparative feedback $\mathcal{F}(a,a^{\prime})=1$ is based on true cost difference $\mu(a)-\mu(a^{\prime})$), then the uncorrupted gradient $g_{t}:=\mathcal{F}(a^{\prime}_{t},a_{t})d\nabla^{2}\mathcal{R}_{t}(a_{t})^{1/2}u_{t}$ should be an unbiased estimate of $\nabla\bar{P}_{t}(a_{t})$, i.e.
@@
We use $\nabla\bar{P}_{t}(a_{t})$ as a shorthand of $\nabla|_{a=a_{t}}\bar{P}_{t}(a)$. We use $P_{t}(a)$ as a shorthand of $P_{t}(a):=\sigma(\mu(a_{t})-\mu(a))$ equivalent to $\mathbb{P}(a_{t}\succ a)$, and $\hat{P}_{t}(a):=\sigma(\mu(a_{t})-\mu(a)+c_{t}(a_{t},a^{\prime}_{t}))$, equivalent to ^$\hat{\mathbb{P}}(a_{t}\succ a)$ under our modelling assumption. The proof is similar to Lemma 2.1 in Flaxman et al. (2005). Using the Law of total expectation, we have
$\mathbb{E}(g_{t}|a_{t})$ $=\mathbb{E}_{u_{t}}[\mathbb{E}(g_{t}|a_{t},u_{t})]$
@@
Step 2: decomposing regret for dueling bandits into regret of decision and feedback error
Core to our proof is a “regret decomposion” lemma below that upper bounds the regret by two parts: regret of sub-optimal decision making under uncertainty, referred as regret of decision (first term in (2)), and feedback error (second term in (2)).
Lemma (Lemma 3 restated).
The $\text{Reg}_{T}$ of Algorithm 1 satisfies:
$\text{Reg}_{T}\leq\underbrace{\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2L^{\mu}L^{\sigma}_{1}R}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Regret\hskip 2.84526ptof\hskip 2.84526ptDecision}}+\underbrace{2\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Feedback\hskip 2.84526ptError}},$
where $C_{0}:=(2\nu+4L^{\sigma}_{1}\kappa+4R^{\sigma}_{2}(L^{\mu})^{2}+L^{\mu}L^{\sigma}_{1}\kappa)/\lambda$, and $a^{*}_{T}:=\frac{1}{T}a_{1}+(1-\frac{1}{T})a^{*}$.
Proof  The cornerstone of the analysis below is to separate the impact of $b_{t}$ from the total regret.
$\text{Reg}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T})\right]+\frac{L^{\mu}L^{\sigma}_{1}\kappa}{\lambda\eta_{\rho}}+2L^{\mu}L^{\sigma}_{1}R$
@@
$+\frac{L^{\mu}L^{\sigma}_{1}\kappa}{\lambda\eta_{\rho}}+2L^{\mu}L^{\sigma}_{1}R$
$\leq 2\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T}))\right\}+\frac{4L^{\sigma}_{1}\kappa+4R^{\sigma}_{2}(L^{\mu})^{2}+L^{\mu}L^{\sigma}_{1}\kappa}{\lambda\eta_{\rho}}\log T+2L^{\mu}L^{\sigma}_{1}R.$
We get the first inequality by using Lemma 16. We get the last inequality by because
$\bar{P}_{t}(a)-P_{t}(a)\leq\frac{L^{\sigma}_{1}\kappa+R^{\sigma}_{2}(L^{\mu})^{2}}{2}\|\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}\|^{2}\leq\frac{L^{\sigma}_{1}\kappa+R^{\sigma}_{2}(L^{\mu})^{2}}{\lambda\eta_{\rho}t},$
@@
$\mathbb{E}\left\{\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T}))\right\}$ $\leq\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T}))\right\}+2\frac{L^{\sigma}_{1}\kappa+R^{\sigma}_{2}(L^{\mu})^{2}}{\lambda\eta_{\rho}}\log T.$
Lemma 16 (Lemma 5 in Kumagai (2017)).
$\text{Reg}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T})\right]+\frac{L^{\mu}L^{\sigma}_{1}\kappa}{\lambda\eta_{\rho}}+2L^{\mu}L^{\sigma}_{1}R.$
@@
$\mathbb{E}\{\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T})\}$ $\leq\mathbb{E}\left[\nabla\hat{P}^{\top}_{t}(a_{t})(a_{t}-a^{*}_{T})-\frac{L^{\sigma}_{1}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[g_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{L^{\sigma}_{1}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[(\hat{g}_{t}+b_{t})^{\top}(a_{t}-a^{*}_{T})-\frac{L^{\sigma}_{1}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{L^{\sigma}_{1}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]+\mathbb{E}\{b^{\top}_{t}(a_{t}-a^{*}_{T})\},$
where the first and second equality is resulted from Lemma 2. Using the definition of $a_{t+1}$ in Algorithm 1, we have
@@
$\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{L^{\sigma}_{1}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$ $=\frac{1}{\eta_{\rho}}\mathbb{E}\left[(\nabla\mathcal{R}_{t}(a_{t+1})-\nabla\mathcal{R}_{t}(a_{t}))^{\top}(a_{t}-a^{*}_{T})-\frac{L^{\sigma}_{1}\alpha\eta_{\rho}}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\frac{1}{\eta_{\rho}}\mathbb{E}\left[D_{\mathcal{R}_{t}}(a^{*}_{T},a_{t})+D_{\mathcal{R}}(a_{t},a_{t+1})-D_{\mathcal{R}_{t}}(a^{*}_{T},a_{t+1})-\frac{L^{\sigma}_{1}\alpha\eta_{\rho}}{4}\|a_{t}-a^{*}_{T}\|^{2}\right].$
@@
Then we can get (15) by using Lemma 17, 18, 19.
@@
Lemma 17 (Lemma 10 in Kumagai (2017)).
@@
$\sum^{T}_{t=1}\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{L^{\sigma}_{1}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]\leq\frac{1}{\eta_{\rho}}\left(\mathcal{R}(a^{*}_{T})-\mathcal{R}(a_{1})+\mathbb{E}\left[\sum^{T}_{t=1}D_{\mathcal{R}^{*}_{t}}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t},\nabla\mathcal{R}_{t}(a_{t}))\right]\right).$
Lemma 18 (Lemma 11 in Kumagai (2017)).